# Spectral selectors on lens spaces and applications to the geometry of the group of contactomorphisms

### v1

@@ 2. The non-linear Maslov index
Before proving Section 2, recall that the cohomological index is subadditive in the following sense ([16, Proposition 3.9 (iv)]): for any two closed subsets $A$ and $B$ of $L_{k}^{2M-1}$ we have
@@
The first inequality and the second one in the case $k=2$ are proved in [16, Proposition 4.15]. Suppose thus that $\{\phi_{t}\}_{t\in[0,1]}$ is negative and $\mu(\{\phi_{t}\}_{t\in[0,t_{0}]})$ is even. By monotonicity of generating functions ([16, Proposition 2.23]), there is a based family of conical generating functions $F_{t}$ for $\{\phi_{t}\}_{t\in[0,1]}$ with $\frac{df_{t}}{dt}\leq 0$. The proof of [16, Proposition 2.23] actually shows that for every $t$ either $\frac{df_{t}}{dt}<0$ or $\frac{df_{t}}{dt}=0$. We can thus assume that either $\{f_{t_{0}}\leq 0\}=\{f_{\underline{t}}\leq 0\}$ or $\{f_{t_{0}}\leq 0\}$ is included in the interior of $\{f_{\underline{t}}\leq 0\}$. In the first case, by continuity of the cohomological index and since the map $s\mapsto\mu(\{\phi_{t}\}_{t\in[0,s]})$ is constant on $(\underline{t},t_{1}]$ we have $\mu(\{\phi_{t}\}_{t\in[0,t_{0}]})=\mu(\{\phi_{t}\}_{t\in[0,t_{1}]})$, which implies the desired inequality
@@
As in the proof of [16, Proposition 4.15] we have
@@
is even (by assumption) and $\operatorname{ind}\big{(}\{f_{0}\leq 0\}\big{)}$ is even (by [16, Remark 3.13]), using (5) we see that $\operatorname{ind}\big{(}\{f_{\underline{t}}\leq-\epsilon\}\big{)}$ is even and so, by subadditivity of the cohomological index,
@@
Since $\operatorname{ind}(C)=\operatorname{ind}\big{(}\Delta(\phi_{\underline{t}})\big{)}$ by [16, Proposition 2.22], using (6) and (7) we thus conclude that
@@
The following Poincaré duality property for the cohomological index is proved in [8, Proposition 4.1.15] for subsets of complex projective space. We adapt here the proof to the case of lens spaces.
@@
Let $\operatorname{ind}(A)=\operatorname{ind}(\mathcal{U}_{A})=2a$. By definition of the cohomological index (cf. [16, Lemma 3.3]) and since Čech cohomology agrees with singular cohomology on open sets, the homomorphism $H^{2a-1}(L_{k}^{2M-1};\mathbb{Z}_{k})\rightarrow H^{2a-1}(\mathcal{U}_{A};\mathbb{Z}_{k})$ induced by the inclusion $\mathcal{U}_{A}\hookrightarrow L_{k}^{2M-1}$ is injective. Since $k$ is prime, the coefficient ring $\mathbb{Z}_{k}$ is a field and so cohomology is the dual of homology, thus the homomorphism $H_{2a-1}(\mathcal{U}_{A};\mathbb{Z}_{k})\rightarrow H_{2a-1}(L_{k}^{2M-1};\mathbb{Z}_{k})$ induced by the inclusion $\mathcal{U}_{A}\hookrightarrow L_{k}^{2M-1}$ is surjective. Consider the commutative square
@@
Since $F_{0}$ is equivalent to the zero function, up to a fibre preserving conical homeomorphism it is equal to a $\mathbb{Z}_{k}$-invariant quadratic form $Q_{0}$. Since $F_{0}$ generates the identity, by [16, Proposition 2.22] the nullity of $Q_{0}$ is equal to $2n$. Let $\iota$ be the dimension of the maximal subspace on which $Q_{0}$ is negative definite. Then
@@
So far we have assumed that $k$ is prime. Suppose now that $k$ is not prime, and let $k^{\prime}$ be the smallest prime that divides $k$. As in [16, Remark 1.4], we define the non-linear Maslov index
@@
For any integer $k\geq 2$ the non-linear Maslov index
@@
    if $k$ is even; in particular, $\mu:\widetilde{\operatorname{Cont}_{0}}(L_{k}^{2n-1},\xi_{0})\rightarrow\mathbb{Z}$ is a quasimorphism.
@@
    Relation with discriminant points: for any contact isotopy $\{\phi_{t}\}_{t\in[0,1]}$ of $(L_{k}^{2n-1},\xi_{0})$ starting at the identity and any subinterval $[t_{0},t_{1}]$ of $[0,1]$, if $\mu(\{\phi_{t}\}_{t\in[0,t_{0}]})\neq\mu(\{\phi_{t}\}_{t\in[0,t_{1}]})$ then there is $\underline{t}\in[t_{0},t_{1}]$ such that $\bar{\phi}_{\underline{t}}$ has discriminant points. Assume that the map $s\mapsto\mu(\{\phi_{t}\}_{t\in[0,s]})$ is constant on $[t_{0},\underline{t})$ and on $(\underline{t},t_{1}]$, and denote by $\Delta_{k^{\prime}}(\phi_{\underline{t}})\subset L_{k^{\prime}}^{2n-1}$ the set of discriminant points at time $\underline{t}$ of the lift of $\{\phi_{t}\}$ to $L_{k^{\prime}}^{2n-1}$, where $k^{\prime}$ is the smallest prime dividing $k$. Then
@@
In the case when $k$ is prime all the properties have been discussed above in Section 2, (1), (2), Section 2, Section 2, Section 2, Section 2, (4), Section 2, Section 2 and Section 2. If $k$ is not prime and $k^{\prime}$ is the smallest prime that divides $k$ then the properties of $\mu:\widetilde{\operatorname{Cont}_{0}}(L_{k^{\prime}}^{2n-1},\xi_{0})\rightarrow\mathbb{Z}$ imply the corresponding properties for the pullback $\mu:\widetilde{\operatorname{Cont}_{0}}(L_{k}^{2n-1},\xi_{0})\rightarrow\mathbb{Z}$ by the natural map $\widetilde{\operatorname{Cont}_{0}}(L_{k}^{2n-1},\xi_{0})\rightarrow\widetilde{\operatorname{Cont}_{0}}(L_{k^{\prime}}^{2n-1},\xi_{0})$. ∎
@@ 3. Spectral selectors
For any $j\in\mathbb{Z}$ we define the $j$-th spectral selector on $\widetilde{\operatorname{Cont}_{0}}(L_{k}^{2n-1},\xi_{0})$ by
@@
$\left\lceil c_{j}\big{(}\widetilde{\phi}\big{)}\right\rceil_{\frac{2\pi}{k}}=\min\left\{\,N\in\frac{2\pi}{k}\cdot\mathbb{Z}\;\ \Big{\lvert}\;\mu(\widetilde{r_{-N}}\cdot\widetilde{\phi})\leq-j\,\right\}\,.$ (10)
@@
We say that a contactomorphism of $(L_{k}^{2n-1},\xi_{0})$ is non-degenerate with respect to $\alpha_{0}$ if all its translated points with respect to $\alpha_{0}$ are non-degenerate. We then have the following lemma.
@@
The set of contactomorphisms of $(L_{k}^{2n-1},\xi_{0})$ contact isotopic to the identity that are non-degenerate with respect to $\alpha_{0}$ is dense in $\operatorname{Cont}_{0}(L_{k}^{2n-1},\xi_{0})$ for the $\mathcal{C}^{1}$-topology.
@@
We first prove the following result. Let $\Lambda_{0}$ be a closed Legendrian submanifold of a contact manifold $\big{(}M,\xi=\ker(\alpha)\big{)}$, and denote by $\mathcal{L}eg(\Lambda_{0})$ its Legendrian isotopy class. Then for any submanifold $N$ of $M$ the set $\mathcal{L}eg_{N}(\Lambda_{0})$ of elements of $\mathcal{L}eg(\Lambda_{0})$ transverse to $N$ is dense in $\mathcal{L}eg(\Lambda_{0})$ for the $\mathcal{C}^{1}$-topology. Indeed, let $\Lambda$ be an element of $\mathcal{L}eg(\Lambda_{0})$ and let $\mathcal{U}(\Lambda)\subset M$ be a Weinstein neighborhood of $\Lambda$. Fix a diffeomorphism $\Psi$ from $\mathcal{U}(\Lambda)$ to an open neighborhood $\mathcal{U}(j^{1}0)$ of the zero section of $J^{1}\Lambda$ such that $\Psi(\Lambda)=j^{1}0$ and $\Psi^{\ast}(dz-\lambda_{\operatorname{can}})=\alpha$. Denote the submanifold $\Psi\big{(}N\cap\mathcal{U}(\Lambda)\big{)}$ of $\mathcal{U}(j^{1}0)$ by $N^{\prime}$. Since the map $j^{1}:\mathcal{C}^{\infty}(\Lambda)\rightarrow\mathcal{L}eg(j^{1}0)$ that associates to a function its 1-jet is a local homeomorphism with respect to the $C^{2}$-topology on $\mathcal{C}^{\infty}(\Lambda)$ and the $\mathcal{C}^{1}$-topology on $\mathcal{L}eg(j^{1}0)$ (cf. for instance [23, Section 3]), for a sufficiently small $\mathcal{C}^{2}$-neighborhood $\mathcal{U}(0)$ of the zero function in $\mathcal{C}^{\infty}(\Lambda)$ the 1-jet of any $f\in\mathcal{U}(0)$ is in $\mathcal{U}(j^{1}0)$ and the map $\mathcal{U}(0)\rightarrow\mathcal{L}eg(\Lambda_{0})$ that sends $f$ to $\Psi^{-1}(j^{1}f)$ is a local homeomorphism. By Thom’s transversality theorem (see for instance [15, Corollary 4.10]), the subset $T^{1}_{N^{\prime}}$ of $\mathcal{U}(0)$ consisting of functions with 1-jet transverse to $N^{\prime}$ is dense in $\mathcal{U}(0)$ for the $\mathcal{C}^{\infty}$-topology, hence also for the $\mathcal{C}^{2}$-topology. Thus for any $\mathcal{C}^{1}$-neighborhood $\mathcal{U}$ of $\Lambda$ in $\mathcal{L}eg(\Lambda_{0})$ there exists $f\in T^{1}_{N^{\prime}}$ such that $\Psi^{-1}(j^{1}f)\in\mathcal{U}$. The Legendrian $\Psi^{-1}(j^{1}f)$ intersects $N$ transversely, and so belongs to $\mathcal{L}eg_{N}(\Lambda_{0})$. This shows that $\mathcal{L}eg_{N}(\Lambda_{0})$ is $\mathcal{C}^{1}$-dense in $\mathcal{L}eg(\Lambda_{0})$.
@@
Using this, we now prove that the set of contactomorphisms of $(L_{k}^{2n-1},\xi_{0})$ contact isotopic to the identity that are non-degenerate with respect to $\alpha_{0}$ is $\mathcal{C}^{1}$-dense in $\operatorname{Cont}_{0}(L_{k}^{2n-1},\xi_{0})$. Consider the product $L_{k}^{2n-1}\times L_{k}^{2n-1}\times\mathbb{R}$ endowed with the contact structure given by the kernel of the contact form $\pi_{2}^{\ast}\alpha_{0}-e^{\theta}\pi_{1}^{\ast}\alpha_{0}$, where $\pi_{1}$ and $\pi_{2}$ denote the projections on the first and second factor respectively and where $\theta$ is the coordinate in $\mathbb{R}$. Consider the closed submanifold $N=\bigcup_{t\in[0,\frac{2\pi}{k}]}\,r_{t}(\Delta\times\{0\})$, where $\{r_{t}\}$ denotes the Reeb flow of $\pi_{2}^{\ast}\alpha_{0}-e^{\theta}\pi_{1}^{\ast}\alpha_{0}$ and $\Delta$ denotes the diagonal. A contactomorphism $\phi$ of $(L_{k}^{2n-1},\xi_{0})$ is non-degenerate with respect to $\alpha_{0}$ if and only if its graph
$\operatorname{gr}_{\alpha_{0}}(\phi)=\big{\{}\big{(}p,\phi(p),g(p)\big{)}\;\big{\lvert}\;p\in L_{k}^{2n-1}\big{\}}\,,$
@@
where $g$ is the conformal factor of $\phi$ with respect to $\alpha_{0}$, is transverse to $N$. In other words, using the notation of the first part of the proof and since $\Delta\times\{0\}=\operatorname{gr}_{\alpha_{0}}(\mathrm{id})$, $\phi$ is non-degenerate with respect to $\alpha_{0}$ if and only if $\operatorname{gr}_{\alpha_{0}}(\phi)\in\mathcal{L}eg_{N}(\operatorname{gr}_{\alpha_{0}}(\mathrm{id}))$. Our result thus follows from the first part of the proof and the fact that the map $\operatorname{gr}_{\alpha_{0}}:\operatorname{Cont}_{0}(L_{k}^{2n-1},\xi_{0})\rightarrow\mathcal{L}eg(\operatorname{gr}_{\alpha_{0}}(\mathrm{id}))$ that associates to a contactomorphism its graph is a local homeomorphism with respect to the $\mathcal{C}^{1}$-topologies. ∎
@@
For any $\widetilde{\phi}\in\widetilde{\operatorname{Cont}_{0}}(L_{k}^{2n-1},\xi_{0})$, $\mathcal{A}(\widetilde{\phi})$ and $\bar{\mathcal{A}}(\widetilde{\phi})$ are closed and nowhere dense in $\mathbb{R}$.
@@
As in [22, Section 5.2 and Lemma 4.4], zero is a regular value of $f$ and so $f^{-1}(0)$ is a submanifold of $L_{k}^{2n(N+1)-1}\times[\frac{1}{2},1]$. Let $p:f^{-1}(0)\rightarrow\big{[}\frac{1}{2},1\big{]}$ be the composition of the inclusion $f^{-1}(0)\hookrightarrow L_{k}^{2n(N+1)-1}\times[\frac{1}{2},1]$ with the projection on the second factor. Then $(x,t)\in f^{-1}(0)$ is a critical point of $p$ (of critical value $t$) if and only if $x$ is a critical point of $f_{t}$ (of critical value zero). Thus the $\mathbb{Z}_{k}$-orbits of translated points of $\bar{\phi}_{1}$ of translation $2\pi(2t-1)$ are in 1–1 correspondence with the critical points of $p$ of critical value $t$. It follows that
@@
and so $\bar{\mathcal{A}}(\widetilde{\phi})$ is closed and nowhere dense. This implies that $\mathcal{A}(\widetilde{\phi})=\bar{\mathcal{A}}(\widetilde{\phi})+\frac{2\pi}{k}\cdot\mathbb{Z}$ is also closed and nowhere dense. ∎
@@
for every $\widetilde{\phi}$ in $\widetilde{\operatorname{Cont}_{0}}(L_{k}^{2n-1},\xi_{0})$. Suppose by contradiction that this is not the case. Since $\bar{\mathcal{A}}(\widetilde{\phi})$ is closed (by Section 3), there is then $\epsilon>0$ such that $[c_{j}(\widetilde{\phi})-\epsilon,c_{j}(\widetilde{\phi})+\epsilon]\subset\mathbb{R}\smallsetminus\bar{\mathcal{A}}(\widetilde{\phi})$. By the first statement of Section 2 (v) we thus have
@@
In particular, $c_{-2n+1}(\widetilde{\mathrm{id}})=c_{0}(\widetilde{\mathrm{id}})=0$.
@@
Suppose that $\widetilde{\phi}\cdot\widetilde{\psi}^{-1}$ is represented by a contact isotopy with Hamiltonian function $H_{t}:L_{k}^{2n-1}\rightarrow\mathbb{R}$ with respect to $\alpha_{0}$. Let $m(t)=\min H_{t}$ and $M(t)=\max H_{t}$. The flows of $m$ and $M$ are respectively $\big{\{}r_{\int^{t}_{0}m}\big{\}}$ and $\big{\{}r_{\int^{t}_{0}M}\big{\}}$, thus
@@
We now show that (11) implies that each $c_{j}$ is continuous with respect to the $\mathcal{C}^{1}$-topology.
Notice first that the Shelukhin–Hofer norm $\nu_{\alpha}:\widetilde{\operatorname{Cont}_{0}}(L_{k}^{2n-1},\xi_{0})\rightarrow\mathbb{R}$, which is defined by

### v2

@@ 2. The non-linear Maslov index
Before proving Section 2, recall that the cohomological index is subadditive in the following sense ([17, Proposition 3.9 (iv)]): for any two closed subsets $A$ and $B$ of $L_{k}^{2M-1}$ we have
@@
The first inequality and the second one in the case $k=2$ are proved in [17, Proposition 4.15]. Suppose thus that $\{\phi_{t}\}_{t\in[0,1]}$ is negative and $\mu(\{\phi_{t}\}_{t\in[0,t_{0}]})$ is even. By monotonicity of generating functions ([17, Proposition 2.23]), there is a based family of conical generating functions $F_{t}$ for $\{\phi_{t}\}_{t\in[0,1]}$ with $\frac{df_{t}}{dt}\leq 0$. The proof of [17, Proposition 2.23] actually shows that for every $t$ either $\frac{df_{t}}{dt}<0$ or $\frac{df_{t}}{dt}=0$. We can thus assume that either $\{f_{t_{0}}\leq 0\}=\{f_{\underline{t}}\leq 0\}$ or $\{f_{t_{0}}\leq 0\}$ is included in the interior of $\{f_{\underline{t}}\leq 0\}$. In the first case, by continuity of the cohomological index and since the map $s\mapsto\mu(\{\phi_{t}\}_{t\in[0,s]})$ is constant on $(\underline{t},t_{1}]$ we have $\mu(\{\phi_{t}\}_{t\in[0,t_{0}]})=\mu(\{\phi_{t}\}_{t\in[0,t_{1}]})$, which implies the desired inequality
@@
As in the proof of [17, Proposition 4.15] we have
@@
is even (by assumption) and $\operatorname{ind}\big{(}\{f_{0}\leq 0\}\big{)}$ is even (by [17, Remark 3.13]), using (5) we see that $\operatorname{ind}\big{(}\{f_{\underline{t}}\leq-\epsilon\}\big{)}$ is even and so, by subadditivity of the cohomological index,
@@
Since $\operatorname{ind}(C)=\operatorname{ind}\big{(}\Delta(\phi_{\underline{t}})\big{)}$ by [17, Proposition 2.22], using (6) and (7) we thus conclude that
@@
The following Poincaré duality property for the cohomological index is proved in [9, Proposition 4.1.15] for subsets of complex projective space. We adapt here the proof to the case of lens spaces.
@@
Let $\operatorname{ind}(A)=\operatorname{ind}(\mathcal{U}_{A})=2a$. By definition of the cohomological index (cf. [17, Lemma 3.3]) and since Čech cohomology agrees with singular cohomology on open sets, the homomorphism $H^{2a-1}(L_{k}^{2M-1};\mathbb{Z}_{k})\rightarrow H^{2a-1}(\mathcal{U}_{A};\mathbb{Z}_{k})$ induced by the inclusion $\mathcal{U}_{A}\hookrightarrow L_{k}^{2M-1}$ is injective. Since $k$ is prime, the coefficient ring $\mathbb{Z}_{k}$ is a field and so cohomology is the dual of homology, thus the homomorphism $H_{2a-1}(\mathcal{U}_{A};\mathbb{Z}_{k})\rightarrow H_{2a-1}(L_{k}^{2M-1};\mathbb{Z}_{k})$ induced by the inclusion $\mathcal{U}_{A}\hookrightarrow L_{k}^{2M-1}$ is surjective. Consider the commutative square
@@
Since $F_{0}$ is equivalent to the zero function, up to a fibre preserving conical homeomorphism it is equal to a $\mathbb{Z}_{k}$-invariant quadratic form $Q_{0}$. Since $F_{0}$ generates the identity, by [17, Proposition 2.22] the nullity of $Q_{0}$ is equal to $2n$. Let $\iota$ be the dimension of the maximal subspace on which $Q_{0}$ is negative definite. Then
@@
So far we have assumed that $k$ is prime. Suppose now that $k$ is not prime, and let $k^{\prime}$ be the smallest prime that divides $k$. As in [17, Remark 1.4], we define the non-linear Maslov index
@@
For any integer $k\geq 2$ and $n$-tuple $\underline{w}=(w_{1},\cdots,w_{n})$ of positive integers relatively prime to $k$, the non-linear Maslov index
$\mu:\widetilde{\operatorname{Cont}_{0}}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}\rightarrow\mathbb{Z}$
@@
    if $k$ is even; in particular, $\mu$ is a quasimorphism.
@@
    Relation with discriminant points: for any contact isotopy $\{\phi_{t}\}_{t\in[0,1]}$ of $\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$ starting at the identity and any subinterval $[t_{0},t_{1}]$ of $[0,1]$, if $\mu(\{\phi_{t}\}_{t\in[0,t_{0}]})\neq\mu(\{\phi_{t}\}_{t\in[0,t_{1}]})$ then there is $\underline{t}\in[t_{0},t_{1}]$ such that $\bar{\phi}_{\underline{t}}$ has discriminant points. Assume that the map $s\mapsto\mu(\{\phi_{t}\}_{t\in[0,s]})$ is constant on $[t_{0},\underline{t})$ and on $(\underline{t},t_{1}]$, and denote by $\Delta_{k^{\prime}}(\phi_{\underline{t}})\subset L_{k^{\prime}}^{2n-1}(\underline{w})$ the set of discriminant points at time $\underline{t}$ of the lift of $\{\phi_{t}\}$ to $L_{k^{\prime}}^{2n-1}(\underline{w})$, where $k^{\prime}$ is the smallest prime dividing $k$. Then
@@
In the case when $k$ is prime all the properties have been discussed above in Section 2, (1), (2), Section 2, Section 2, Section 2, Section 2, (4), Section 2, Section 2 and Section 2. If $k$ is not prime and $k^{\prime}$ is the smallest prime that divides $k$ then the properties of $\mu:\widetilde{\operatorname{Cont}_{0}}\big{(}L_{k^{\prime}}^{2n-1}(\underline{w}),\xi_{0}\big{)}\rightarrow\mathbb{Z}$ imply the corresponding properties for the pullback $\mu:\widetilde{\operatorname{Cont}_{0}}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}\rightarrow\mathbb{Z}$ by the natural map $\widetilde{\operatorname{Cont}_{0}}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}\rightarrow\widetilde{\operatorname{Cont}_{0}}\big{(}L_{k^{\prime}}^{2n-1}(\underline{w}),\xi_{0}\big{)}$. ∎
@@ 3. Spectral selectors
For any $j\in\mathbb{Z}$ we define the $j$-th spectral selector on $\widetilde{\operatorname{Cont}_{0}}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$ by
@@
$\left\lceil c_{j}\big{(}\widetilde{\phi}\big{)}\right\rceil_{T_{\underline{w}}}=\min\left\{\,N\in T_{\underline{w}}\cdot\mathbb{Z}\;\ \Big{\lvert}\;\mu(\widetilde{r_{-N}}\cdot\widetilde{\phi})\leq-j\,\right\}\,.$ (10)
@@
We say that a contactomorphism of $\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$ is non-degenerate with respect to $\alpha_{0}$ if all its translated points with respect to $\alpha_{0}$ are non-degenerate for all their translations. We then have the following lemma.
@@
The set of contactomorphisms of $\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$ contact isotopic to the identity that are non-degenerate with respect to $\alpha_{0}$ is dense in $\operatorname{Cont}_{0}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$ for the $\mathcal{C}^{1}$-topology.
@@
We first prove the following result. Let $\Lambda_{0}$ be a closed Legendrian submanifold of a contact manifold $\big{(}M,\xi=\ker(\alpha)\big{)}$, and denote by $\mathcal{L}eg(\Lambda_{0})$ its Legendrian isotopy class. Then for any collection $N=\{N_{l}\,,\,l\in\mathbb{Z}\}$ of submanifolds of $M$ the set $\mathcal{L}eg_{N}(\Lambda_{0})$ of elements of $\mathcal{L}eg(\Lambda_{0})$ transverse to $N_{l}$ for all $l$ is dense in $\mathcal{L}eg(\Lambda_{0})$ for the $\mathcal{C}^{1}$-topology. Indeed, let $\Lambda$ be an element of $\mathcal{L}eg(\Lambda_{0})$ and let $\mathcal{U}(\Lambda)\subset M$ be a Weinstein neighborhood of $\Lambda$. Fix a diffeomorphism $\Psi$ from $\mathcal{U}(\Lambda)$ to an open neighborhood $\mathcal{U}(j^{1}0)$ of the zero section of $J^{1}\Lambda$ such that $\Psi(\Lambda)=j^{1}0$ and $\Psi^{\ast}(dz-\lambda_{\operatorname{can}})=\alpha$. Denote for each $l$ by $N_{l}^{\prime}$ the submanifold $\Psi\big{(}N_{l}\cap\mathcal{U}(\Lambda)\big{)}$ of $\mathcal{U}(j^{1}0)$, and let $N^{\prime}=\{N^{\prime}_{l}\,,\,l\in\mathbb{Z}\}$. Since the map $j^{1}:\mathcal{C}^{\infty}(\Lambda)\rightarrow\mathcal{L}eg(j^{1}0)$ that associates to a function its 1-jet is a local homeomorphism with respect to the $C^{2}$-topology on $\mathcal{C}^{\infty}(\Lambda)$ and the $\mathcal{C}^{1}$-topology on $\mathcal{L}eg(j^{1}0)$ (cf. for instance [24, Section 3]), for a sufficiently small $\mathcal{C}^{2}$-neighborhood $\mathcal{U}(0)$ of the zero function in $\mathcal{C}^{\infty}(\Lambda)$ the 1-jet of any $f\in\mathcal{U}(0)$ is in $\mathcal{U}(j^{1}0)$ and the map $\mathcal{U}(0)\rightarrow\mathcal{L}eg(\Lambda_{0})$ that sends $f$ to $\Psi^{-1}(j^{1}f)$ is a local homeomorphism. By Thom’s transversality theorem (see for instance [16, Corollary 4.10]), the subset $T^{1}_{N^{\prime}}$ of $\mathcal{U}(0)$ consisting of functions with 1-jet transverse to $N_{l}^{\prime}$ for all $l$ is dense in $\mathcal{U}(0)$ for the $\mathcal{C}^{\infty}$-topology, hence also for the $\mathcal{C}^{2}$-topology. Thus for any $\mathcal{C}^{1}$-neighborhood $\mathcal{U}$ of $\Lambda$ in $\mathcal{L}eg(\Lambda_{0})$ there exists $f\in T^{1}_{N^{\prime}}$ such that $\Psi^{-1}(j^{1}f)\in\mathcal{U}$. The Legendrian $\Psi^{-1}(j^{1}f)$ intersects each $N_{l}$ transversely, and so belongs to $\mathcal{L}eg_{N}(\Lambda_{0})$. This shows that $\mathcal{L}eg_{N}(\Lambda_{0})$ is $\mathcal{C}^{1}$-dense in $\mathcal{L}eg(\Lambda_{0})$.
@@
Using this, we now prove that the set of contactomorphisms of $\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$ contact isotopic to the identity that are non-degenerate with respect to $\alpha_{0}$ is $\mathcal{C}^{1}$-dense in $\operatorname{Cont}_{0}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$. Consider the contact product of $\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$, i.e. the product $L_{k}^{2n-1}(\underline{w})\times L_{k}^{2n-1}(\underline{w})\times\mathbb{R}$ endowed with the contact structure given by the kernel of the contact form $\pi_{2}^{\ast}\alpha_{0}-e^{\theta}\pi_{1}^{\ast}\alpha_{0}$, where $\pi_{1}$ and $\pi_{2}$ denote the projections on the first and second factor respectively and where $\theta$ is the coordinate in $\mathbb{R}$. Denote by $\operatorname{gr}_{\alpha_{0}}(\phi)$ the graph with respect to $\alpha_{0}$ of a contactomorphism $\phi$ of $\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$, i.e. the Legendrian submanifold
$\operatorname{gr}_{\alpha_{0}}(\phi)=\big{\{}\big{(}p,\phi(p),g(p)\big{)}\;\big{\lvert}\;p\in L_{k}^{2n-1}(\underline{w})\big{\}}$
@@
of the contact product, where $g$ is the conformal factor of $\phi$ with respect to $\alpha_{0}$. Denote by $\{R_{t}\}$ the Reeb flow of $\pi_{2}^{\ast}\alpha_{0}-e^{\theta}\pi_{1}^{\ast}\alpha_{0}$. Let $0=t_{0}<\cdots<t_{m}=2\pi$ be a decomposition of the time interval $[0,2\pi]$ such that, for some $\epsilon>0$, $N_{l}:=\bigcup_{t\in(t_{l}-\epsilon,t_{l+1}+\epsilon)}\,R_{t}\big{(}\operatorname{gr}_{\alpha_{0}}(\mathrm{id})\big{)}$ is a submanifold of $L_{k}^{2n-1}(\underline{w})\times L_{k}^{2n-1}(\underline{w})\times\mathbb{R}$ for every $l\in\{0,\cdots,m-1\}$. A contactomorphism $\phi$ of $\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$ is non-degenerate with respect to $\alpha_{0}$ if and only if $\operatorname{gr}_{\alpha_{0}}(\phi)$ is transverse to $N_{l}$ for all $l$. In other words, using the notation of the first part of the proof, $\phi$ is non-degenerate with respect to $\alpha_{0}$ if and only if $\operatorname{gr}_{\alpha_{0}}(\phi)\in\mathcal{L}eg_{N}(\operatorname{gr}_{\alpha_{0}}(\mathrm{id}))$, where $N=\{N_{0},\cdots,N_{m-1}\}$. Our result thus follows from the first part of the proof and the fact that the map
$\operatorname{gr}_{\alpha_{0}}:\operatorname{Cont}_{0}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}\rightarrow\mathcal{L}eg(\operatorname{gr}_{\alpha_{0}}(\mathrm{id}))$
that associates to a contactomorphism its graph is a local homeomorphism with respect to the $\mathcal{C}^{1}$-topologies. ∎
@@
For any element $\widetilde{\phi}$ of $\widetilde{\operatorname{Cont}_{0}}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$, the sets $\mathcal{A}(\widetilde{\phi})$ and $\bar{\mathcal{A}}(\widetilde{\phi})$ are closed and nowhere dense in $\mathbb{R}$.
@@
As in [23, Section 5.2 and Lemma 4.4], zero is a regular value of $f$ and so $f^{-1}(0)$ is a submanifold of $L_{k}^{2n(N+1)-1}\times[\frac{1}{2},1]$. Let $p:f^{-1}(0)\rightarrow\big{[}\frac{1}{2},1\big{]}$ be the composition of the inclusion $f^{-1}(0)\hookrightarrow L_{k}^{2n(N+1)-1}\times[\frac{1}{2},1]$ with the projection on the second factor. Then $(x,t)\in f^{-1}(0)$ is a critical point of $p$ (of critical value $t$) if and only if $x$ is a critical point of $f_{t}$ (of critical value zero). Thus the $\mathbb{Z}_{k}$-orbits of translated points of $\bar{\phi}_{1}$ of translation $2\pi(2t-1)$ are in 1–1 correspondence with the critical points of $p$ of critical value $t$. It follows that
@@
and so $\bar{\mathcal{A}}(\widetilde{\phi})$ is closed and nowhere dense. Since
$\bar{\mathcal{A}}(\widetilde{\phi})+T_{\underline{w}}\cdot\mathbb{Z}\subseteq\mathcal{A}(\widetilde{\phi})\subseteq\bar{\mathcal{A}}(\widetilde{\phi})+\frac{2\pi}{k}\cdot\mathbb{Z}\,,$
we deduce that $\mathcal{A}(\widetilde{\phi})$ is also closed and nowhere dense. ∎
@@
for every $\widetilde{\phi}$ in $\widetilde{\operatorname{Cont}_{0}}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$. Suppose by contradiction that this is not the case. Since $\bar{\mathcal{A}}(\widetilde{\phi})$ is closed (by Section 3), there is then $\epsilon>0$ such that $[c_{j}(\widetilde{\phi})-\epsilon,c_{j}(\widetilde{\phi})+\epsilon]\subset\mathbb{R}\smallsetminus\bar{\mathcal{A}}(\widetilde{\phi})$. By the first statement of Section 2 (v) we thus have
@@
In particular, $c_{0}(\widetilde{\mathrm{id}})=0$.
@@
Suppose that $\widetilde{\phi}\cdot\widetilde{\psi}^{-1}$ is represented by a contact isotopy with Hamiltonian function $H_{t}:L_{k}^{2n-1}(\underline{w})\rightarrow\mathbb{R}$ with respect to $\alpha_{0}$. Let $m(t)=\min H_{t}$ and $M(t)=\max H_{t}$. The flows of $m$ and $M$ are respectively $\big{\{}r_{\int^{t}_{0}m}\big{\}}$ and $\big{\{}r_{\int^{t}_{0}M}\big{\}}$, thus
@@
We now show that (11) implies that each $c_{j}$ is continuous with respect to the $\mathcal{C}^{1}$-topology. Notice first that the Shelukhin–Hofer norm $\nu_{\alpha}:\widetilde{\operatorname{Cont}_{0}}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}\rightarrow\mathbb{R}$, which is defined by